40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Feb 9 22:43:29 2021, rana, Noise Hunting, Laser, "North" current driver having slow current fluctuations PXL_20210210_063043141.jpg
    Reply  Wed Feb 10 09:52:45 2021, shruti, Noise Hunting, Laser, "North" current driver having slow current fluctuations 
    Reply  Mon Feb 15 21:25:00 2021, rana, Noise Hunting, Laser, "North" current driver having slow current fluctuations Planex-Beatnote.pdfPlanex-Beatnote-ASD.pdfplotPhaseMeter.tar.gz
    Reply  Thu Feb 18 12:30:09 2021, aaron, Noise Hunting, Laser, "North" current driver having slow current fluctuations 
Message ID: 2636     Entry time: Tue Feb 9 22:43:29 2021     Reply to this: 2637   2645   2651
Author: rana 
Type: Noise Hunting 
Category: Laser 
Subject: "North" current driver having slow current fluctuations 

I spent this evening trying to find the source of the slow laser frequency fluctuations I see on the Moku.

The fluctuations are ~150 MHz peak-peak, whereas we had very steady performance before (I can't read the .mat file Shruti sent me; not sure if it got corrupted during transfer or if Moke records in a weird .mat format - I couldn't load it in matlab or python).

Looking at the current outptu of the driver I see a relaxation oscillation with a ~8 s period. This seems to correspond to the observed frequency fluctuations. I bet there's something wrong in the driver.

In the attached scope image, the yellow trace is the DC output of the power meter, the purple trace is the "total current output" of the current driver, and the blue trace is the DC output of the beatnote PD.

 

Also, the new labeling is a little confusing for me. Its a little hard to follow which path is north/south when the fibers go in and out of the isolators, etc. Probabl have to add more labels.

 

The new fiber mounts are too high. We should put them on an adjustable post for now, but get some other posts machined so that these give a 4" beam height. Check the Thorlabs CAD drawing.

 

We really need to scan the temperature of each laser and record the power to see where the mode-hop free zones are, or else we're flying blind.

Also, it would be good to take the TEC outputs on the back of the TEC controllers and route them into the DAQ so that we have these values all the time. Same with the diode currents. Are the DAQ channels being recorded or is the Cymac down?

Attachment 1: PXL_20210210_063043141.jpg  427 kB  Uploaded Wed Feb 10 08:20:35 2021  | Hide | Hide all
PXL_20210210_063043141.jpg
ELOG V3.1.3-